OpenStudy (anonymous):

Simplify by combining like terms. -8m+11n+6+5m-9n-10+y

OpenStudy (anonymous):

lets group them together first \[-8m+5m+11n-9n+6-10+y\]

OpenStudy (anonymous):

now combine since \(-8+5=-3\) we get \(-3m\) since \(11-9=2\) we have \(2n\) also \(6-10=-4\) and \(y\) has no match that gives \[-3m+2n-4+y\]

OpenStudy (anonymous):

Thank you @satellite73

OpenStudy (anonymous):

yw, with practice you do this with you eyeballs, no need to rearrange

OpenStudy (anonymous):

btw kudos to your teacher (or book) for writing "simplify by combining like terms" and not just "simplify"